OpenStudy (anonymous):

AC=3/4CB. Find C??? Help

OpenStudy (anonymous):

Move the RHS to the LHS and then factor the LHS. Do you get the following for a result?\[\frac{1}{4} (4 A-3 B) C=0\]

OpenStudy (anonymous):

dont understand

OpenStudy (anonymous):

\[\frac{1}{4} (4 A - 3 B) C \text{ = } 0\]Divide each side by the coefficient of C.\[C \text{ = }\frac{ 0}{\frac{1}{4} (4 A - 3 B)} \]\[C = 0 \]